OpenStudy (hulkart):

Calculate the coefficient of x2y6 in the expansion of (x + y)8. 15 12 56 28

OpenStudy (hulkart):

@gaos

OpenStudy (anonymous):

\[\binom{8}{5}\] or \[\binom{8}{3}\] same number

OpenStudy (anonymous):

Shouldn't it be \[\left(\begin{matrix}8 \\ 2\end{matrix}\right)\] or \[\left(\begin{matrix}8 \\ 6\end{matrix}\right)\] ?

OpenStudy (loser66):

yes, it is @SemantcSpider. @Satelite73 made mistake.

OpenStudy (anonymous):

@Loser66 it's @SemanticSpider and @satellite73

OpenStudy (anonymous):

and yes @satellite73 made a mistake